• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 08:33:03 on July23,2003

    Entry number 106306

    keyword=Septum Training

    We put the septum set current to 375 A, in preperation to do more training.